Quintron a local communication company makes parts for the Delta 4 rocket, the microchips used in the communication system must be within .001 milligrams of the population mean.  Determine the minimum required sample size to construct a 90% confidence interval for the true mean weight of microchips. Assume the population standard deviation is .01 milligrams
Answer #1

Solution :

Given that,

Population standard deviation = = 0.01

Margin of error = E = 0.001

At 90% confidence level the z is,

= 1 - 90%

= 1 - 0.90 = 0.10

/2 = 0.05

Z/2 = Z0.05 = 1.645

sample size = n = [Z/2* / E] 2

n = [1.645 * 0.01/ 0.001]2

n = 270.60

Sample size = n = 271